Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4c)
direnzo - 23:07 Thursday 28 August 2025 (67587) Print this report
Operator Report - Afternoon shift

At the start of the shift, the interferometer just recovered from an unlock likely due to intense wind, and was in Science mode for about 30 minutes. Another unlock occurred at 13:41:19 UTC, GPS 1440423697. After a series of unlocks for the strong wind (50+ km/h) before reaching LOW_NOISE_3 state, at 14:35 UTC the interferometer has been set to Commissioning mode, moving up the scheduled activities. LOW_NOISE_3 recovered at 18:57 UTC, after the calibration activities.

Below, the list of performed activities:

  • 14:35 UTC - INJ, realignment of the input beam (De Rossi, Casanueva): #67588. Interferometer set to LOCKED_ARMS_IR. Process concluded at 15:30 UTC
  • 15:30 - Napoli team to NEB for microphone and tiltmeter calibration. Operation concluded at 17:00 UTC
  • 15:30 - Started the calibration measurements with ITF unlocked:
    • 15:35 - Started Measurement of the TF between CAL and Sc channels, and of PCal sensing delays. Concluded at 15:48 UTC

    • 16:20 - Measurement of actuators response for PR and BS mirrors. Issue with DET_MAIN shutter closed when setting CALIBRATED_PRWI. Restored by putting DET_MAIN back to INIT, and restarting the calibration procedure. Concluded at 17:17 UTC

    • 17:20 - Measurement of actuators response for NI,WI mirrors. Concluded at 18:40 UTC

  • 15:48 - TCS, Cifaldi check of the powers. Operation competed at 16:18 UTC

  • 19:13 - CALI, Verkindt: Hrec bias updated. 

  • The interfermeter unlocked again at 19:30:51 UTC, GPS 1440444669, when a hailstorm hit the site, and, given that we were already 30 miutes past the end of the Commissioning slot, we decided to postpone the Scattered Light calibration activity by van Haevermaet. The interferometer was set beck to PREPARE_ASCIENCE mode (and AUTOSCIENCE_ON)

Other events to report:

  • Sudden data loss at 16:00:09 UTC. See attached figure
  • At 19:59 UTC, after various failed locking attempts and ITF mode set to BAD_WEATHER. The strong wind even tore down the Belgian flag!
  • At 20:24, ITF reached again LOW_NOISE_3 and set back to AUTO_SCIENCE

Guard tours (time in UTC)

  • 18:03 - 18:39
  • 21:00 - ...
Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×